A boat leaves X and travels east for 10km. It then sails 10km south to Y. Find the distance and bearing from X from Y
Answer by meraj(5)

Let the distance XY=10km be the base of the triangle and distance
YZ=10km be another base of the triangle.
Now,
XZ^2=XY^2+YZ^2
XZ^2=10^2+10^2
XZ^2=100+100
XZ^2=200
XZ=14.14213562373095km